# Hidden Markov Models and the Bayes Filter in Categorical Probability

### 26 Jan 2024 v1

@@ 1 Introduction
footnotetext: Acknowledgments. We thank Dario Stein for suggesting Proposition 3.2 and Matthew B. Smith for help with the forward-backward algorithm. This research was funded in part by the Austrian Science Fund (FWF) [doi:10.55776/P35992, doi:10.55776/P33122, doi:10.55776/P34129]. For open access purposes, the authors have applied a CC BY public copyright license to any author accepted manuscript version arising from this submission. AK further acknowledges funding of the Austrian Academy of Sciences (ÖAW) through the DOC scholarship 26547. YW’s research is further supported by the University of Innsbruck Early Stage Funding Program.
@@ 1.4 Detailed outline
For a hidden Markov kernel, we may think of $\mathrm{BF}_{t}^{\sharp}$ as the Bayes filter in a form that deterministically outputs the observer’s posterior distribution at time $t$ as a function of the observations. In 3.20, we show that in every representable Markov category, the deterministic Bayes filters form a Markov chain of their own, in the sense that the probability distribution in $PX_{n}$ exclusively depends on the attained distribution in $PX_{n-1}$.
@@ 2.1 Markov chains
    In $\mathsf{FinSetMulti}$, a Markov chain is again described by (2.5), but now with a possibility distribution $f_{0}\colon X_{0}\to\mathopen{}\mathclose{{}\left\{0,1}\right\}$ and $\mathopen{}\mathclose{{}\left\{0,1}\right\}$-valued matrices $f_{t}\colon X_{t-1}\times X_{t}\to\mathopen{}\mathclose{{}\left\{0,1}\right\}$ such that for every $x_{t}\in X_{t}$ there is at least one $x_{t-1}\in X_{t-1}$ with $f_{t}\mathopen{}\mathclose{{}\left(x_{t}\,|\,x_{t-1}}\right)=1$. Equivalently, $f_{0}$ is a subset of $X_{0}$ consisting of those elements considered as possible initial states, while $f_{t}$ is a subset of $X_{t-1}\times X_{t}$ telling us for a given state at time $t-1$ which states at time $t$ are possible. By (2.2), the joint $p\colon I\to\bigotimes_{t=0}^{n}X_{t}$ is then a subset of $\prod_{t=0}^{n}X_{t}$ consisting of those tuples $\mathopen{}\mathclose{{}\left(x_{0},\ldots,x_{n}}\right)$ which can occur as possible trajectories of the Markov chain.
@@ 3.1 The Bayes filter
Virgo’s work adopts a perspective slightly different from ours. Fixing an observation space $Y$, he forms a category $\mathsf{Generator}\mathopen{}\mathclose{{}\left(Y}\right)$ having generalized hidden Markov models (in the above sense) with observation space $Y$ as objects and morphisms those maps between the state spaces which preserve the structure [45, Section 2.1]. Assuming that the Markov category in question is strongly representable, Virgo then extends Bayesian filtering to a functor out of the category $\mathsf{Generator}\mathopen{}\mathclose{{}\left(Y}\right)$. This can be seen as functorially converting a hidden Markov model into a machine that updates Bayesian priors. It is further shown that this Bayesian filtering functor is a right adjoint [45, Theorem 2.6]. While Virgo’s hidden Markov models are partly more restrictive (time-homogeneous) and partly more general than ours, his work is complementary to ours insofar as it is concerned with different aspects of the Bayes filter. For instance, Virgo does not prove any version of Proposition 3.2, focusing instead on the adjunction of [45, Theorem 2.6] and its consequences (such as constructing conjugate priors).
@@ 3.3.3 The instantiated Bayes filter in $\mathsf{FinSetMulti}$
Next, we compute the Bayes filter in a ‘‘possibilistic’’ case, that of $\mathsf{FinSetMulti}$. Recall from A.22(e) that a conditional $f_{|X}$ of an arbitrary morphism $f\colon A\to X\otimes Y$ is given simply by
@@
$q_{t}\mathopen{}\mathclose{{}\left(x_{t}}\right)=\sum_{x_{t-1}}f_{t}\mathopen{}\mathclose{{}\left(x_{t}\,|\,x_{t-1}}\right)\cdot\mathrm{IBF}_{t-1}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[t-1}\right]}}\right)\mathopen{}\mathclose{{}\left(x_{t-1}}\right).$
This now means that $x_{t}$ is considered possible if and only if $x_{t-1}$ is possible according to our previous estimate of the hidden state $\mathrm{IBF}_{t-1}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[t-1}\right]}}\right)$ and $x_{t}$ is possible according to the transition kernel $f_{t}$. The update step is given simply by
$\mathrm{IBF}_{t}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[t}\right]}}\right)\mathopen{}\mathclose{{}\left(x_{t}}\right)=g_{t}\mathopen{}\mathclose{{}\left(y_{t}\,|\,x_{t}}\right)\cdot q_{t}\mathopen{}\mathclose{{}\left(x_{t}}\right).$
@@
As far as we know, this has not been studied before in the context of automata theory. We call this version of the Bayes filter the possibilistic filter. [^17] We omit mention of Bayes here, since this version of the Bayes filter no longer uses Bayesian updating in the usual sense.
@@ 3.4 The Bayes filter recursion as a single string diagram
Since the recursion of 3.12 is easy to expand, we immediately arrive at the following.
@@
Consequently, as with Remark 3.4 and Remark 3.8, the results of this section also apply to $\mathcal{C}_{W}$. This immediately shows that 3.13 holds even in case all morphisms involved have an additional dependence on some parameter object $W$. This amounts to every box in our diagrams, including the $\mathrm{u}_{i}$, connecting to a copy of the $W$ input as in (2.6).
@@ 3.5 The Bayes filter as a Markov chain
Applying 3.19, this evaluates further to
@@
where the second equation uses the recursion formula of 3.12. By the fact that $\mathrm{BF}_{t-1}^{\sharp}$ is the deterministic counterpart of $\mathrm{BF}_{t}$, we can also write this as
@@
where the second equation holds by 3.16 and the definition of $h_{t}$. The claim now follows straightforwardly by the induction hypothesis. ∎
@@ A.4 Conditionals and conditioning
    See also [43, Section 6.5] for a more geometric construction of conditionals in a closely related category of extended Gaussians.

### 25 Feb 2024 v2

@@ 1 Introduction
footnotetext: Acknowledgments. We thank Dario Stein for suggesting Proposition 3.2, Matthew B. Smith for help with the forward-backward algorithm, and Ramon van Handel for helpful discussion about non-linear filtering. This research was funded in part by the Austrian Science Fund (FWF) [doi:10.55776/P35992, doi:10.55776/P33122, doi:10.55776/P34129]. For open access purposes, the authors have applied a CC BY public copyright license to any author accepted manuscript version arising from this submission. AK further acknowledges funding of the Austrian Academy of Sciences (ÖAW) through the DOC scholarship 26547. YW’s research is further supported by the University of Innsbruck Early Stage Funding Program.
@@ 1.3 Overview of the paper
As a companion work to the paper, we have also prepared an implementation of our formalism in the programming language C++, written by Matvey Soloviev and the last named author, at [42]. We briefly discuss this implementation in Section 1.5.
@@ 1.4 Detailed outline
For a hidden Markov kernel, we may think of $\mathrm{BF}_{t}^{\sharp}$ as the Bayes filter in a form that deterministically outputs the observer’s posterior distribution at time $t$ as a function of the observations. In 3.21, we show that in every representable Markov category, the deterministic Bayes filters form a Markov chain of their own, in the sense that the probability distribution in $PX_{n}$ exclusively depends on the attained distribution in $PX_{n-1}$.
@@
1.5 Implementation
In the context of computing, the abstract nature of Markov categories allows any algorithm expressed in terms of string diagrams to be implemented in a reusable way, as it can be instantiated both in a probabilistic and possibilistic context, or generally in any Markov category that may be of interest to the programmer. As a demonstration, Matvey Soloviev and the last named author have implemented a proof-of-concept codebase [42] that formalizes Markov categories as C++ abstract classes [45], which are a mechanism that enables writing programs that operate on any data structure for which a particular set of methods has been implemented.
Based on this, we then provide an implementation of the instantiated Bayes filter (Section 3.2) that mirrors the string-diagrammatic formulas, together with implementations of the categories $\mathsf{FinStoch}$ and $\mathsf{FinSetMulti}$, in which the filter implementation can be used with no change to its code. To add support for an additional category, the programmer needs only to implement a representation of objects and morphisms in that category, together with methods that implement the Markov category structure as well as conditioning. [^4] Depending on the category, it may be advantageous to implement some frequently-used operations which can be derived from the basic ones directly, exploiting additional structure in the category for increased performance. Our implementation demonstrates an example of this for the projection morphisms $A_{1}\otimes\ldots\otimes A_{n}\rightarrow A_{i}$, which can be obtained as a tensor product of $n-1$ deletion morphisms with an identity, but has a faster implementation in $\mathsf{FinStoch}$ when the dimensions of the $A_{i}$ are known.
@@ 2.1 Markov chains
    In $\mathsf{FinSetMulti}$, a Markov chain is again described by (2.5), but now with a possibility distribution $f_{0}\colon X_{0}\to\mathopen{}\mathclose{{}\left\{0,1}\right\}$ and $\mathopen{}\mathclose{{}\left\{0,1}\right\}$-valued matrices $f_{t}\colon X_{t-1}\times X_{t}\to\mathopen{}\mathclose{{}\left\{0,1}\right\}$ such that for every $x_{t}\in X_{t}$ there is at least one $x_{t+1}\in X_{t+1}$ with $f_{t+1}\mathopen{}\mathclose{{}\left(x_{t+1}\,|\,x_{t}}\right)=1$. Equivalently, $f_{0}$ is a subset of $X_{0}$ consisting of those elements considered as possible initial states, while $f_{t}$ is a subset of $X_{t-1}\times X_{t}$ telling us for a given state at time $t-1$ which states at time $t$ are possible. By (2.2), the joint $p\colon I\to\bigotimes_{t=0}^{n}X_{t}$ is then a subset of $\prod_{t=0}^{n}X_{t}$ consisting of those tuples $\mathopen{}\mathclose{{}\left(x_{0},\ldots,x_{n}}\right)$ which can occur as possible trajectories of the Markov chain.
@@ 3.1 The Bayes filter
Virgo’s work adopts a perspective slightly different from ours. Fixing an observation space $Y$, he forms a category $\mathsf{Generator}\mathopen{}\mathclose{{}\left(Y}\right)$ having generalized hidden Markov models (in the above sense) with observation space $Y$ as objects and morphisms those maps between the state spaces which preserve the structure [47, Section 2.1]. Assuming that the Markov category in question is strongly representable, Virgo then extends Bayesian filtering to a functor out of the category $\mathsf{Generator}\mathopen{}\mathclose{{}\left(Y}\right)$. This can be seen as functorially converting a hidden Markov model into a machine that updates Bayesian priors. It is further shown that this Bayesian filtering functor is a right adjoint [47, Theorem 2.6]. While Virgo’s hidden Markov models are partly more restrictive (time-homogeneous) and partly more general than ours, his work is complementary to ours insofar as it is concerned with different aspects of the Bayes filter. For instance, Virgo does not prove any version of Proposition 3.2, focusing instead on the adjunction of [47, Theorem 2.6] and its consequences (such as constructing conjugate priors).
@@ 3.3.3 The instantiated Bayes filter in $\mathsf{FinSetMulti}$
Next, we compute the Bayes filter in a ‘‘possibilistic’’ case, that of $\mathsf{FinSetMulti}$. As far as we know, this has not been studied before in the context of automata theory. We call this version of the Bayes filter the possibilistic filter. [^18] We omit mention of Bayes here, since this version of the Bayes filter no longer uses Bayesian updating in the usual sense.
Recall from A.22(e) that a conditional $f_{|X}$ of an arbitrary morphism $f\colon A\to X\otimes Y$ is given simply by
@@
$q_{t}\mathopen{}\mathclose{{}\left(x_{t}}\right)=\sum_{x_{t-1}}f_{t}\mathopen{}\mathclose{{}\left(x_{t}\,|\,x_{t-1}}\right)\cdot\mathrm{IBF}_{t-1}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[t-1}\right]}}\right)\mathopen{}\mathclose{{}\left(x_{t-1}}\right).$ (3.9)
This now means that $x_{t}$ is considered possible if and only if $x_{t-1}$ is possible according to our previous estimate of the hidden state $\mathrm{IBF}_{t-1}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[t-1}\right]}}\right)$ and $x_{t}$ is possible according to the transition kernel $f_{t}$. Choosing conditionals as in A.22(e) and using the general prescription shows that the update step is given simply by
$\mathrm{IBF}_{t}\mathopen{}\mathclose{{}\left(\mathbf{y}_{\mathopen{}\mathclose{{}\left[t}\right]}}\right)\mathopen{}\mathclose{{}\left(x_{t}}\right)=g_{t}\mathopen{}\mathclose{{}\left(y_{t}\,|\,x_{t}}\right)\cdot q_{t}\mathopen{}\mathclose{{}\left(x_{t}}\right).$ (3.10)
@@
Example 3.8.
Let us work through a simple example of a hidden Markov model in $\mathsf{FinSetMulti}$ and the possibilistic filter. Consider first the time-homogeneous Markov chain—or equivalently as a nondeterministic automaton—on a three-element set $X=\{a,b,c\}$ with possible transitions like this:
{tikzpicture}
Our transition kernels $f_{t}$ are all the same for $t>0$ by time-homogeneity, and we thus omit the index $t$. The transition kernel $f$ can then also be written in matrix form as [^19] This means that $f(x^{\prime}|x)=1$ if it is possible to reach $x^{\prime}$ from $x$ in one step. This matrix representation is the transpose of transition matrices as conventionally used for Markov chains in classical probability theory.
$f=(f(x^{\prime}|x))_{x,x^{\prime}\in\{a,b,c\}}=\begin{pmatrix}1&0&0\\ 1&0&0\\ 0&1&1\end{pmatrix}.$
Suppose that our chain starts at $a$. This is represented by the morphism $f_{0}:I=\{e\}\to X$, where
$f_{0}(a|e)=1\qquad\text{and}\qquad f_{0}(b|e)=f_{0}(c|e)=0,$
or in matrix notation
$(f_{0}(x|e))_{x\in\{a,b,c\}}=\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}.$
Let us consider observation morphisms $g_{t}:\{a,b,c\}\to\{\bot,\top\}$ that are also independent of $t$ and given by the matrix
$g=(g(y|i))_{y\in\{\bot,\top\},i\in\{a,b,c\}}=\begin{pmatrix}1&1&0\\ 0&1&1\end{pmatrix}.$
This means that the hidden state $a$ is always observed as $\bot$, the hidden state $c$ as $\top$, and the hidden state $b$ as either. Given the underlying Markov chain, it is clear that the possible sequences of observations all looks as follows: a sequence of $\bot$’s as long as the hidden state is $a$, possibly followed by an arbitrary sequence of $\bot$ and $\top$ while it is $b$, possibly followed by only $\top$’s if the chain ends up in its terminal state $c$.
Now, let’s compute the instantiated Bayes filter by doing the prediction and update steps. For time $t=0$, the prediction coincides with the initial distribution: $q_{0}=f_{0}$. We already know that $y_{0}=\bot$ is the only possible observation, for which the initial update step gives
$\mathrm{IBF}_{0}(\bot)=\begin{pmatrix}g(\bot|a)\,q_{0}(a)\\ g(\bot|b)\,q_{0}(b)\\ g(\bot|c)\,q_{0}(c)\end{pmatrix}=\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}.$
This is unsurprising, as it simply tells us what we already knew, namely that the initial state must have been $a$ with certainty. Note that the other instance $\mathrm{IBF}_{0}(\top)$ depends on the choice of conditional, since $y_{0}=\top$ is not possible under the initial state $a$. Using the choice of conditionals that leads to (3.10), we get $\mathrm{IBF}_{0}(\top)=\mathrm{IBF}_{0}(\bot)$, in line with the fact that the initial observation does not tell us anything new about the hidden state.
Suppose now that our sequence of observations up to $t=2$ is
$\mathbf{y}_{[2]}=(\bot,\bot,\top).$
At time $t=1$, the chain can a priori be in state $a$ or $b$. Indeed using the $\mathsf{FinSetMulti}$-arithmetic $1+1=1$ in the prediction step (3.9) gives
$q_{1}=f\;\;\mathrm{IBF}_{0}(\bot)=\begin{pmatrix}1&0&0\\ 1&0&0\\ 0&1&1\end{pmatrix}\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}=\begin{pmatrix}1\\ 1\\ 0\end{pmatrix}.$
Applying the update step (3.10) produces
$\mathrm{IBF}_{1}(\bot,\bot)=\begin{pmatrix}g(\bot|a)\,q_{1}(a)\\ g(\bot|b)\,q_{1}(b)\\ g(\bot|c)\,q_{1}(c)\end{pmatrix}=\begin{pmatrix}1\\ 1\\ 0\end{pmatrix}.$
So with our second observation having been $y_{1}=\bot$, we cannot distinguish between the hidden states $a$ and $b$, but we can be sure that the chain is not in state $c$.
Now at time $t=2$, and before we take the observation $y_{2}=\top$ into account, the chain can be in any of the three states: the prediction step (3.9) gives
$q_{2}=f\;\;\mathrm{IBF}_{1}(\bot,\bot)=\begin{pmatrix}1&0&0\\ 1&0&0\\ 0&1&1\end{pmatrix}\begin{pmatrix}1\\ 1\\ 0\end{pmatrix}=\begin{pmatrix}1\\ 1\\ 1\end{pmatrix}.$
At this point, the third observation $y_{2}=\top$ finally gives us nontrivial information: using the update step (3.10), we find
$\mathrm{IBF}_{2}(\bot,\bot,\bot)=\begin{pmatrix}1\\ 1\\ 0\end{pmatrix},$
which again is intuitively clear as the observed value $y_{2}=\top$ has excluded the hidden state $a$.
@@ 3.4 The Bayes filter recursion as a single string diagram
Since the recursion of 3.13 is easy to expand, we immediately arrive at the following.
@@
Consequently, as with Remark 3.4 and Remark 3.9, the results of this section also apply to $\mathcal{C}_{W}$. This immediately shows that 3.14 holds even in case all morphisms involved have an additional dependence on some parameter object $W$. This amounts to every box in our diagrams, including the $\mathrm{u}_{i}$, connecting to a copy of the $W$ input as in (2.6).
@@ 3.5 The Bayes filter as a Markov chain
Applying 3.20, this evaluates further to
@@
where the second equation uses the recursion formula of 3.13. By the fact that $\mathrm{BF}_{t-1}^{\sharp}$ is the deterministic counterpart of $\mathrm{BF}_{t}$, we can also write this as
@@
where the second equation holds by 3.17 and the definition of $h_{t}$. The claim now follows straightforwardly by the induction hypothesis. ∎
@@
Example 3.23.
In $\mathsf{FinSetMulti}$, each distribution object $PX_{t}$ is the set of nonempty subsets of $X_{t}$. Indeed evaluating (3.14) in $\mathsf{FinSetMulti}$ shows that the filter update
$\mathrm{u}_{t}^{\sharp}\>:\>PX_{t-1}\otimes Y_{t}\longrightarrow PX_{t}$
is the map which takes a subset of $X_{t-1}$ and an observation $y_{t}$, applies the transition $f_{t}$ to determine all states $x_{t}$ consistent with the given set at time $t-1$ and then intersects with the set of states consistent with the observed value $y_{t}$. This is effectively what we already did in 3.8 to compute the possibilistic filter.
Continuing on from that example, let us see what the filter process and its Markov property amount to there. At $t=0$, we have seen that the filter deterministically outputs
$\mathrm{IBF}_{0}(\bot)=\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}.$
At time $t=1$, the possible filter outputs are
$\mathrm{IBF}_{1}(\bot,\bot)=\begin{pmatrix}1\\ 1\\ 0\end{pmatrix},\qquad\mathrm{IBF}_{1}(\bot,\top)=\begin{pmatrix}0\\ 1\\ 0\end{pmatrix},$
depending on the value of the second observation. At time $t=2$, we similarly get one of
$\mathrm{IBF}_{2}(\bot,\bot,\bot)=\begin{pmatrix}1\\ 1\\ 0\end{pmatrix},\qquad\mathrm{IBF}_{2}(\bot,\bot,\top)=\begin{pmatrix}0\\ 1\\ 1\end{pmatrix},\qquad\mathrm{IBF}_{2}(\bot,\top,\top)=\begin{pmatrix}0\\ 0\\ 1\end{pmatrix}.$
Continuing like this, a simple induction argument shows that the filter process is given by the following possibilistic Markov chain, or nondeterministic automaton:
{tikzpicture}
Here, the labels indicate which new observations trigger which transition. Note that $\bot$ does not appear on any transition out of the bottom two states, since when we are at $b$ or $c$ with certainty, then we must be at $c$ in the next step and will necessarily observe $\top$.
@@ A.4 Conditionals and conditioning
    See also [44, Section 6.5] for a more geometric construction of conditionals in a closely related category of extended Gaussians.